# The Race between Technology and Woman:

### v8

@@ A.2.1 Substitution parameters
where the degree of substitution between ICT capital and the four types of labor is governed by a single parameter $\varsigma$. We suppress the sector subscript $n$ for notational simplicity but allow all parameters to vary across sectors. Cost minimization implies that the marginal rate of technical substitution of labor for capital equals the relative price of labor to capital for all types of labor. This condition leads to the following first-difference equations:
@@
Manufacturing –0.698 0.287 0.602 0.782 9.689 90.545 15.856
(0.337) (0.047) (0.043) (0.056) [0.002] [0.000] [0.000]
Utilities –0.904 0.227 0.427 0.657 19.944 6.049 25.385
(0.289) (0.083) (0.057) (0.059) [0.000] [0.014] [0.000]
Services –0.415 0.342 0.599 0.871 26.064 21.328 74.775
@@
Notes: Columns 1 to 4 report the estimates of the substitution parameters with standard errors clustered at the country level in parentheses. Columns 5 to 7 report Wald statistics under the stated hypotheses with p-values in square brackets.
A.2.2 Nesting structure
@@
where the two substitution parameters ($\varsigma_{fh}$, $\varsigma_{fu}$) are less than one, and the two share parameters ($\vartheta_{fh}$, $\vartheta_{fu}$) lie between zero and one. This specification relaxes the assumption that male and female labor are perfect substitutes yet still imposes the assumption that male and female labor are equally substitutable with ICT capital for each skill type. This assumption implies that technological change embodied in new equipment has no effect on the gender wage gap. Both the assumption and the implication are inconsistent with the data, as shown above. Therefore, there are only two alternatives. One is our preferred specification (28). The other is the following specification:
@@ A.2.3 Error terms
A.2.4 Weights in the aggregate elasticities
@@
$\pi_{\ell_{f}\ell_{g},n}^{fh}$ =$=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{1},\ell_{g}\in\mathcal{L}_{1}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{1},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{mh}$ =$=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{mh},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{2},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{1},\ell_{g}=\ell_{mh}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{1},\ell_{g}=\ell_{mh}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{2},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{fu}$ =$=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{fu},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{3},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{2},\ell_{g}=\ell_{fu}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{2},\ell_{g}=\ell_{fu}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{3},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{mu}$ =$=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{mu},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{4},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{3},\ell_{g}=\ell_{mu}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{3},\ell_{g}=\ell_{mu}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{4},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{o}$ =$=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)\lambda_{k_{o}n}&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{4}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)-\pi_{\ell_{f}\ell_{g},n}^{c}&\text{if }\ell_{f}=k_{o}\\ \frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\pi_{\ell_{f}\ell_{g},n}^{c}&\text{if }\ell_{g}=k_{o},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{c}$ =$=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)\left(1-\lambda_{k_{o}n}\right)&\text{if }\ell_{g}\in\mathcal{L}_{4}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\lambda_{\ell_{g}n}&\text{if }\ell_{g}=k_{o},\end{cases}$
@@
A.2.5 Magnitude of the effects of technological change
We show that the effects of technological progress embodied in ICT capital tend to increase as the differences in the substitution parameters in the production function increase and the income share of ICT capital relative to the income shares of other factors increases. For this purpose, we assume here that the aggregate production function is of the nested CES form (28) (i.e., $N=1$) and that the capital–skill–gender complementarity hypothesis holds (i.e., $\sigma_{fh}<\sigma_{mh}<\sigma_{fu}<\sigma_{mu}$). The effects of technological change embodied in ICT capital on the relative wages and income shares of the four types of labor can be obtained by setting $d\ln q_{o}=d\ln A_{n}=0$ in equations (23) and (24), respectively.

### v9

@@ E.1 Substitution parameters
where the degree of substitution between ICCT capital and the four types of labor is governed by a single parameter $\varsigma$. We suppress the sector subscript $n$ for notational simplicity but allow all parameters to vary across sectors. Cost minimization implies that the marginal rate of technical substitution of labor for capital equals the relative price of labor to capital for all types of labor. This condition leads to the following first-difference equations:
@@
Manufacturing –0.689 0.288 0.600 0.776 10.540 108.465 14.488
(0.319) (0.046) (0.043) (0.054) [0.001] [0.000] [0.000]
Utilities –0.962 0.184 0.404 0.631 15.300 6.140 23.543
(0.362) (0.104) (0.063) (0.058) [0.000] [0.013] [0.000]
Services –0.413 0.340 0.598 0.861 25.515 22.783 61.220
@@
Notes: The first to fourth columns report the estimates of the substitution parameters with standard errors clustered at the country level in parentheses. The fifth to seventh columns report Wald statistics under the stated hypotheses with p-values in square brackets.
E.2 Nesting structure
@@
where the two substitution parameters ($\varsigma_{fh}$, $\varsigma_{fu}$) are less than one, and the two share parameters ($\vartheta_{fh}$, $\vartheta_{fu}$) lie between zero and one. This specification relaxes the assumption that male and female labor are perfect substitutes yet still imposes the assumption that male and female labor are equally substitutable with ICCT capital for each skill type. This assumption implies that technological change embodied in new equipment has no effect on the gender wage gap. Both the assumption and the implication are inconsistent with the data, as shown above. Therefore, there are only two alternatives. One is our preferred specification (24). The other is the following specification:
@@ E.3 Error terms
E.4 Weights in the aggregate elasticities
@@
$\pi_{\ell_{f}\ell_{g},n}^{fh}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{1},\ell_{g}\in\mathcal{L}_{1}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{1},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{mh}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{mh},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{1}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mh}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{2},\ell_{g}\in\mathcal{L}_{1}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{1},\ell_{g}=\ell_{mh}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{1},\ell_{g}=\ell_{mh}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{2},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{fu}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{fu},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{2}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{fu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{3},\ell_{g}\in\mathcal{L}_{2}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{2},\ell_{g}=\ell_{fu}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{2},\ell_{g}=\ell_{fu}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{3},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{mu}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}=\ell_{mu},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{3}}\lambda_{\ell_{h}n}}\right)\left(\frac{\lambda_{\ell_{mu}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{4},\ell_{g}\in\mathcal{L}_{3}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)+\left(\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\in\mathcal{L}_{3},\ell_{g}=\ell_{mu}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(1-\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)&\text{if }\ell_{f}\notin\mathcal{L}_{3},\ell_{g}=\ell_{mu}\\ 0&\text{if }\ell_{g}\notin\mathcal{L}_{4},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{o}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)\lambda_{k_{o}n}&\text{if }\ell_{f},\ell_{g}\in\mathcal{L}_{4}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)-\pi_{\ell_{f}\ell_{g},n}^{c}&\text{if }\ell_{f}=k_{o}\\ \frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\pi_{\ell_{f}\ell_{g},n}^{c}&\text{if }\ell_{g}=k_{o},\end{cases}$
@@
$\pi_{\ell_{f}\ell_{g},n}^{c}$ $=\begin{cases}\left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\left(\frac{\lambda_{\ell_{g}n}}{\sum_{\ell_{h}\in\mathcal{L}_{4}}\lambda_{\ell_{h}n}}\right)\left(1-\lambda_{k_{o}n}\right)&\text{if }\ell_{g}\in\mathcal{L}_{4}\\ \left(\frac{\zeta_{n}\lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}-\frac{\zeta_{n}\lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\right)\lambda_{\ell_{g}n}&\text{if }\ell_{g}=k_{o},\end{cases}$
@@
E.5 Magnitude of the effects of technological change
We show that the effects of technological progress embodied in ICCT capital tend to increase as the differences in the substitution parameters in the production function increase and the income share of ICCT capital relative to the income shares of other factors increases. For this purpose, we assume here that the aggregate production function is of the nested CES form (24) (i.e., $N=1$) and that the capital–skill–gender complementarity hypothesis holds (i.e., $\sigma_{fh}<\sigma_{mh}<\sigma_{fu}<\sigma_{mu}$). The effects of technological change embodied in ICCT capital on the relative wages and income shares of the four types of labor can be obtained by setting $d\ln q_{o}=d\ln A_{n}=0$ in equations (A1) and (A2), respectively.